HW 2 6 Saved Help Save & Exit Submit 10 Check my work High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costing system in which departmental predetermined overhead rates are used to apply manufacturing overhead cost to jobs. The predetermined overhead rate in the Molding Department is based on machine-hours, and the rate in the Painting Department is based on direct labor hours. At the beginning of the year, the company provided the following estimates: points Department Molding Painting Direct labor-hours Machine-hours Fixed manufacturing overhead cost Variable nanufacturing overhead per machine-hour Variable manufacturing overhead per direct labor-hour 56, 200 39,000 $234,900 $517,040 31,000 81,000 Book $ 2.60 -$4.60 Print Job 205 was started on August 1 and completed on August 10. The companys cost records show the following information concerning oferencesthe job: MoldingPainting Direct labor-hours Machine-hours Direct materials Direct labor cost 127 68 $ 946 $1,120 $ 720 980 80 380 Required 1. Compute the predetermined overhead rates used in the Molding Department and the Painting Department. 2. Compute the total overhead cost applied to Job 205. 3-a. What would be the total manufacturing cost recorded for Job 205? 3-b. If the job contained 30 units, what would be the unit product cost?
Answer #1
1 Predetermined overhead rate
Molding Department
Fixed manufacturing overhead cost (234,900/81,000)= 2.9
Variable manufacturing overhead rate per machine hour 2.6
Predetermined overhead rate 5.5 per MH's
Painting Department
Fixed manufacturing overhead cost (517040/56200)= 9.2
Variable manufacturing overhead rate per machine hour 4.6
Predetermined overhead rate 13.8 per DLH
2) total overhead cost applied ot job 205
Molding painting total
Molding (380*5.5) 2090 2090
Painting (127*13.8) 1753 1753
total 2090 1753 3843
total overhead cost applied to job 205    3843
3-a) total manufacturing cost
Molding painting total
Direct materials 946 1120 2066
Direct labor cost 720 980 1700
overhead cost applied 2090 1753 3843
total 3756 3853 7609
total manufacturing cost $7,609
3-b) unit product cost = 7609/30
253.63